# Dynamical Quantum Phase Transition and Thermal Equilibrium in the Lattice Thirring Model

### v1

@@ Abstract
Using tensor network methods, we simulate the real-time evolution of the lattice Thirring model quenched out of equilibrium in both the critical and massive phases, and study the appearance of dynamical quantum phase transitions, as non-analyticities in the Loschmidt rate. Whereas the presence of a dynamical quantum phase transition in the model does not correspond to quenches across the critical line of the equilibrium phase diagram at zero temperature, we identify a threshold in the energy density of the initial state, necessary for a dynamical quantum phase transition to be present. Moreover, in the case of the gapped quench Hamiltonian, we unveil a connection of this threshold to a transition between different regions in the finite temperature phase diagram.
@@ I Introduction
Out-of-equilibrium dynamics is one of the most challenging problems in the study of quantum many-body systems. This funnels the interest towards finding universal behaviours that allow a more comprehensive understanding of the dynamics. A paradigmatic scenario to investigate these questions is that of a quantum quench [1], in which a system is evolved with a Hamiltonian of interest, after having been initialized in a certain (often pure) state. In the last decade, an active direction of research in this context has been the investigation of dynamical quantum phase transitions (DQPTs) [2, 3], which appear as zeroes in the Loschmidt echo or return probability, i.e. the probability that the system is found in the initial state during the evolution.
@@
Signatures of DQPT have been observed experimentally in trapped ion quantum simulators [31] and superconducting qubits [32]. While initial studies suggested that DQPTs occur when the parameter of the Hamiltonian is quenched across an equilibrium phase boundary [2, 3, 24], it was later realized that this is not the general case [33, 17, 34, 35, 36], and the nature of DQPTs is more complex, and far from being completely understood [29, 37].
@@
The study of out-of-equilibrium physics for quantum field theories (QFTs), defined in the continuum, is highly non-trivial. Equilibrium properties of such models are usually addressed by the path integral formalism and Monte Carlo sampling, which however encounters a sign problem in certain scenarios. This limitation has motivated an investigation of applicability of TNS methods to the realm of QFT. In particular, gauge theories in 1+1D have been systematically addressed using TNS methods (see e.g. [38, 39, 40] and references therein). Other QFTs and conformal field theories (CFTs) have also been studied in the same manner [41, 42, 43, 44]. These investigations have focused on the equilibrium properties of the field theories. Real-time simulations are more scarce [45, 46, 47, 48], and studying the real time dynamics of a continuum QFT with these methods remains an open problem.
@@
The most straightforward way to study QFT problems with TNS is to use the discretized version of the model Hamiltonian, and apply standard TNS methods to find their equilibrium or dynamical properties. Such results depend on different lattice discretization schemes, but by repeating the procedure at different lattice spacings and performing the proper renormalization, it is possible to recover the continuum limit (see e.g. [49, 45]). In this work, we focus on one such spin chain, corresponding to the discretization of the 1+1 dimensional Thirring model. The resulting spin model is a specialized XXZ Hamiltonian coupled to both staggered and homogeneous magnetic fields. The ground state of the model, as determined in previous work [41] using MPS, exhibits a critical and a gapped phase, separated by a Berezinskii-Kosterlitz-Thoules (BKT) phase transition (see Fig. 1). For vanishing bare fermion mass, the model is critical for all couplings.
@@
In this work, we use uniform MPS to study the phenomenology of DQPTs in the spin model for the discretized Thirring model. We simulate quenches in both phases from initial states with very different properties. We show that a necessary condition for a DQPT to occur within the finite time window that our numerical method can access is that the energy density of the initial state is above a certain threshold, independent of the phases of the initial and quenched Hamiltonian. However, in the case of the quench to the gapped phase we also identify a connection of DQPTs to the equilibrium phase diagram, but at finite temperature. Namely, at the temperature corresponding to the identified energy density threshold, the properties of the thermal equilibrium state undergo a substantial change, with the vanishing of a string correlator that is non-zero in the ground state. We further show that the times at which DQPTs occur as a function of the energy density of the initial state show a structure reminiscent of the complex zeros of the Loschmidt amplitude [50, 51], and several branches of DQPTs that can be explored from different initial states.
@@
The rest of the paper is structured as follows. In Sec. II we introduce the spin model that corresponds to the discretized Thirring Hamiltonian, and present the MPS formulation and the tools and algorithms we have used. Second, we review the equilibrium phase structure of the Thirring model from the previous research in Sec. III. In Sec. IV, we discuss the DQPT concept and the use of MPS methods to study them in the model of interest. Our results for DQPTs in different quenches, and the relation of their appearance with the finite temperature phase diagram are discussed in Sec. V. Finally, in Sec. VI we summarize our findings and discuss potential further investigations.
@@ II.1 The Thirring model as a spin model
In this work, we study the massive Thirring model in 1+1 dimensions. The action is described by
$\begin{split}S_{\mathrm{Th}}[\psi,\bar{\psi}]\ =&\int d^{2}x[\bar{\psi}i\gamma^{\mu}\partial_{\mu}\psi\ -m\,\bar{\psi}\psi\\ &-\frac{g}{2}\left(\bar{\psi}\gamma_{\mu}\psi\right)\left(\bar{\psi}\gamma^{\mu}\psi\right)]\,,\end{split}$ (1)
@@
where $m$ is the fermion mass and $g$ is the four-fermion coupling. To facilitate the use of MPS, quantization of the classical theory described by Eq. (1) is performed through the canonical method in the Hamiltonian formalism. One subtlety in this formalism is the inclusion of effects of the anomaly in the four-fermion operator. These effects can be easily investigated through the examination of the path-integral measure in the Lagrangian formulation [52]. In the canonical procedure, they need to be accounted for via a non-local definition of the currents in the four-fermion operator [53, 54].
@@
The model can be discretized on a one-dimensional spatial lattice using the staggered regularization. For convenience of carrying out numerical computations in this work, the fermionic degrees of freedom in this Hamiltonian are mapped onto spin operators through the Jordan-Wigner (JW) transformation. Details of the above procedure can be found in Ref. [41]. Implementation of this strategy turns the Hamiltonian operator of the continuum Thirring model, $H_{{\mathrm{Th}}}$, into that of the XXZ spin chain coupled to both uniform and staggered magnetic fields, up to a scaling factor,
$H(\tilde{m}_{0}a,\Delta)=-\frac{1}{2}\sum_{n}^{N-2}\left(S_{n}^{+}S_{n+1}^{-}+S_{n+1}^{+}S_{n}^{-}\right)+a\tilde{m}_{0}\sum_{n}^{N-1}\left(-1\right)^{n}\left(S_{n}^{z}+\frac{1}{2}\right)+\Delta(g)\sum_{n}^{N-1}\left(S_{n}^{z}+\frac{1}{2}\right)\ \left(S_{n+1}^{z}+\frac{1}{2}\right)\,,$ (2)
@@
where $a$ is the lattice spacing, $N$ is the total number of lattice sites, $S_{n}^{\pm}=S_{n}^{x}\pm iS_{n}^{y}$ and $S_{n}^{z}$ are the spin matrices ($S^{i}_{n}=\sigma^{i}/2$ with $\sigma^{i}$ being the Pauli matrices) at the $n{-}$th site, and $[S^{i}_{n},S^{j}_{m}]_{n\not=m}=0$. Here, $\tilde{m}_{0}=m_{0}/\nu(g)$ with $m_{0}$ being the bare counterpart of the mass parameter, $m$, in Eq. (1). The functions $\nu(g)$ and $\Delta(g)$ are the lattice version of wavefunction renormalization and the four-fermion coupling [55],
$\begin{split}\nu(g)=&\left(\frac{\pi-g}{\pi}\right)/\sin\left(\frac{\pi-g}{2}\right)\,,\\ &\Delta(g)=\cos\left(\frac{\pi-g}{2}\right)\,.\end{split}$ (3)
@@
Notice that the physical Hamiltonian is obtained by rescaling the dimensionless one in Eq. (2) by a factor $\nu(g)/a$. Throughout the rest of the paper, we will work with the lattice Hamiltonian  (2) and correspondingly define dimensionless time and inverse temperature parameters, which involve rescaling the physical (dimensionful) magnitudes by $\nu(g)/a$. Additionally, we restrict our consideration to the sector with total spin (magnetization) zero [41] and find the ground state by adding the penalty term (See appendix  A.)
@@ II.2 Matrix product state
which can then be approximated as a matrix product state (MPS) with bond dimension $D$,
@@
We search for DQPTs in the real-time evolution of Eq. (2). For this purpose, it is desirable to perform calculations directly in the thermodynamic limit. This can be accomplished by employing the uniform MPS (uMPS) technique, which allows for the description of the state of an infinite spin chain with translation invariance using one bulk tensor and two appropriate boundary tensors [56, 57]. The uniform MPS can be represented as
@@
where $A$ is the unit cell of the uMPS. Due to the staggered term in Eq. (2), the system is translationally invariant with a period of two. Thus, we choose the tensor $A$ to represent two contiguous sites of the chain, and its physical dimension becomes $d=4$.
III Equilibrium phase structure of the 1+1 dimensional Thirring model
In this section, we review our study of the zero-temperature phase structure of the Thirring model using MPS [41]. An analysis of renormalization group (RG) equations of the dual sine-Gordon theory [58, 59, 60] points to the existence of two phases in the Thirring model. The massless theory is a conformal field theory at all couplings $\Delta(g)$. RG flows in the mass-coupling plane reveal that the line $m=0$ is a fixed line under RG transformations. However, it is stable/unstable for couplings $\Delta(g)$ below/above certain $\Delta(g_{\star})_{m=0}=-\pi/2$. At $m>0$, the non-zero fermion mass is an irrelevant operator for $\Delta(g)<\Delta(g_{\star})_{m}$, but becomes relevant when $\Delta(g)>\Delta(g_{\star})_{m}$, with the transition coupling $\Delta(g_{\star})_{m}$ being mass-dependent. This gives rise to two equilibrium phases:
  • $\Delta(g)<\Delta(g_{\star})_{m}$ – gapless (critical) phase,
  • $\Delta(g)>\Delta(g_{\star})_{m}$ – gapped (massive) phase,
@@ III Equilibrium phase structure of the 1+1 dimensional Thirring model
$\begin{split}C_{\rm string}(x)=&\frac{1}{N_{x}}\sum_{n}\langle S^{z}(n)S^{z}(n+1)\cdots\\ &\cdots S^{z}(n+x-1)S^{z}(n+x)\rangle,\end{split}$ (8)
which is a string of $S^{z}$ operators. We note that Ref. [41] erroneously states this string correlator as the fermion-antifermion one, containing $\langle S^{+}(n)S^{z}(n+1)\ldots S^{z}(n+x-1)S^{-}(n+x)\rangle$ terms. At large enough distances, the correlator $C_{\rm string}(x)$ exhibits a power-law decay in the gapless phase, given by $C_{\rm string}(x)\propto\beta x^{-\alpha}$, and a power-exponential decay in the gapped phase, $C_{\rm string}(x)\propto Bx^{-\eta}A^{x}$. The crucial distinction that makes the string correlator particularly useful for determining $\Delta(g_{\star})_{m}$ is that it decays to a constant, $C$, in the gapped phase. In practice, one can analyze both phases based on the power-exponential fitting ansatz:
@@
with $A=1$ found in fits in the gapless phase. Thus, the phase transition when going from the gapless into the gapped phase is manifested by $A$ dropping below 1 and, in the case of the string correlator, by $C>0$. The onset of the non-zero constant is numerically more robust than the smooth drop of $A$ below 1, accounting for the practical usefulness of the string correlator. As we will see below, this is true also at non-zero temperatures and it will allow us to identify relations between thermal phases of the Thirring model and DQPTs.
As a summary of our findings for the zero-temperature phase structure, we show Fig. 1. The color-coded regions correspond to the equilibrium phase on the coupling-mass plane through the correlator analyses from our previous work [41]. In particular, we fitted Eq. (9) and extracted values of the constant $C$. Since the results depend, in general, on the fitting interval, we adopted a systematic procedure to assess the systematic uncertainty of $C$ (dominating with respect to other uncertainties, coming from the finite bond dimension and the finite volume). For details of this procedure, we refer again to Ref. [41]. Depending on the value of $C$, we identified three situations. With $C<0.001$ or $C>0.01$, one unambiguously concludes either the critical phase (purple region) or the gapped phase (orange region), respectively. If $C\in[0.001,0.01]$ (grey region), the systematic uncertainty does not allow us to draw conclusions about a (non)vanishing value of $C$. This region of ambiguous values of $C$ is indicated by grey shading and the BKT transition must occur there. We note the onset a nonzero value of $C$ is very smooth at small fermion masses, while it becomes rather sharp with an increasing mass. At very small masses, the implied transition point is consistent with the RG flow analysis ($\Delta(g)\approx-\pi/2$) and at large masses, it moves towards more negative values.
Figure 1: Zero-temperature phase diagram of the Thirring model in the coupling-mass plane. The color-coded areas on the plot depict the equilibrium phase through correlator analyses [41]. The blue region is the critical region, where the constant $C$ (defined in Eq. (9)) is less than $0.001$; the orange region is the gapped phase region, where $C>0.01$; the grey region is undetermined, where the corresponding $C\in[0.001,0.01]$. The black point marks the location where we observe the DQPT, starting from the ground state of the Hamiltonian $H_{1}(m_{1}a,\Delta_{1})$ and quenching to $H_{2}(m_{2}a=0.1,\Delta_{2}=0.9)$.

### v2

@@ Abstract
Using tensor network methods, we simulate the real-time evolution of the lattice Thirring model quenched out of equilibrium in both the critical and massive phases and study the appearance of dynamical quantum phase transitions, as nonanalyticities in the Loschmidt rate. Although the presence of a dynamical quantum phase transition in the model does not correspond to quenches across the critical line of the equilibrium phase diagram at zero temperature, we identify a threshold in the energy density of the initial state, necessary for a dynamical quantum phase transition to be present. Moreover, in the case of the gapped quench Hamiltonian, we unveil a connection of this threshold to a transition between different regions in the finite-temperature phase diagram.
@@ I Introduction
Out-of-equilibrium dynamics is one of the most challenging problems in the study of quantum many-body systems. This funnels the interest towards finding universal behaviors that allow a more comprehensive understanding of the dynamics. A paradigmatic scenario to investigate these questions is that of a quantum quench [1], in which a system is evolved with a Hamiltonian of interest, after having been initialized in a certain (often pure) state. Since the mid-2010s, an active direction of research in this context has been the investigation of dynamical quantum phase transitions (DQPTs) [2, 3], which appear as zeros in the Loschmidt echo or return probability, i.e., the probability that the system is found in the initial state during the evolution.
@@
Signatures of DQPT have been observed experimentally in trapped ion quantum simulators [31] and superconducting qubits [32]. While initial studies suggested that DQPTs occur when the parameter of the Hamiltonian is quenched across an equilibrium phase boundary [2, 3, 24], it was later realized that this is not the general case [33, 17, 34, 35, 36], and the nature of DQPTs is more complex and far from being completely understood [29, 37].
@@
The study of out-of-equilibrium physics for quantum field theories (QFTs), defined in the continuum, is highly nontrivial. Equilibrium properties of such models are usually addressed by the path integral formalism and Monte Carlo sampling, which, however, encounters a sign problem in certain scenarios. This limitation has motivated an investigation of applicability of TNS methods to the realm of QFT. In particular, gauge theories in 1+1 dimensions have been systematically addressed using TNS methods (see, e.g., Refs. [38, 39, 40] and references therein). Other QFTs and conformal field theories have also been studied in the same manner [41, 42, 43, 44]. These investigations have focused on the equilibrium properties of the field theories. Real-time simulations are more scarce [45, 46, 47, 48, 49], and studying the real-time dynamics of a continuum QFT with these methods remains an open problem.
@@
The most straightforward way to study QFT problems with TNS is to use the discretized version of the model Hamiltonian and apply standard TNS methods to find their equilibrium or dynamical properties. Such results depend on different lattice discretization schemes, but by repeating the procedure at different lattice spacings and performing the proper renormalization, it is possible to recover the continuum limit (see, e.g., Refs.  [50, 45]). In this work, we focus on one such spin chain, corresponding to the discretization of the 1+1- dimensional Thirring model. The resulting spin model is a specialized XXZ Hamiltonian coupled to both staggered and homogeneous magnetic fields. The ground state of the model, as determined in previous work [41] using MPS, exhibits a critical and a gapped phase, separated by a Berezinskii-Kosterlitz-Thoules (BKT) phase transition (see Fig. 1). For vanishing bare fermion mass, the model is critical for all couplings.
Figure 1: Zero-temperature phase diagram of the Thirring model in the coupling-mass plane. The color-coded areas on the plot depict the equilibrium phase through correlator analyses [41]. The blue region is the critical region, where the constant $C$ [defined in Eq. (9)] is less than $0.001$; the orange region is the gapped phase region, where $C>0.01$; the gray region is undetermined, where the corresponding $C\in[0.001,0.01]$. The black point marks the location where we observe the DQPT, starting from the ground state of the Hamiltonian $H_{1}(m_{1}a,\Delta_{1})$ and quenching to $H_{2}(m_{2}a=0.1,\Delta_{2}=0.9)$.
@@
In this work, we use uniform MPS to study the phenomenology of DQPTs in the spin model for the discretized Thirring model. We simulate quenches in both phases from initial states with very different properties. We show that a necessary condition for a DQPT to occur within the finite-time window that our numerical method can access is that the energy density of the initial state is above a certain threshold, independent of the phases of the initial and quenched Hamiltonian. However, in the case of the quench to the gapped phase we also identify a connection of DQPTs to the equilibrium phase diagram but at finite temperature. Namely, at the temperature corresponding to the identified energy density threshold, the properties of the thermal equilibrium state undergo a substantial change, with the vanishing of a string correlator that is nonzero in the ground state. We further show that the times at which DQPTs occur as a function of the energy density of the initial state show a structure reminiscent of the complex zeros of the Loschmidt amplitude [51, 52] and several branches of DQPTs that can be explored from different initial states.
@@
The rest of the paper is structured as follows. In Sec. II we introduce the spin model that corresponds to the discretized Thirring Hamiltonian and present the MPS formulation and the tools and algorithms we have used. Second, we review the equilibrium phase structure of the Thirring model from the previous research in Sec. III. In Sec. IV, we discuss the DQPT concept and the use of MPS methods to study them in the model of interest. Our results for DQPTs in different quenches, and the relation of their appearance with the finite temperature phase diagram are discussed in Sec. V. Finally, in Sec. VI we summarize our findings and discuss potential further investigations.
@@ II.1 The Thirring model as a spin model
In this work we focus on the spin chain formulation corresponding to the lattice discretization of the massive Thirring model in 1+1 dimensions. The action of the original field theory is described by
$S_{\text{Th}}\left[\psi,\bar{\psi}\right]=$ $\int d^{2}x\Big{[}\bar{\psi}i\gamma^{\mu}\partial_{\mu}\psi-m\bar{\psi}\psi$
$-\frac{g}{2}(\bar{\psi}\gamma_{\mu}\psi)(\bar{\psi}\gamma^{\mu}\psi)\Big{]},$ (1)
@@
where $m$ is the fermion mass and $g$ is the four-fermion coupling. To facilitate the use of MPS, quantization of the classical theory described by Eq. (1) is performed through the canonical method in the Hamiltonian formalism. One subtlety in this formalism is the inclusion of effects of the anomaly in the four-fermion operator. These effects can be easily investigated through the examination of the path-integral measure in the Lagrangian formulation [53]. In the canonical procedure, they need to be accounted for via a nonlocal definition of the currents in the four-fermion operator [54, 55].
@@
The model can be discretized on a one-dimensional spatial lattice using the staggered regularization. For convenience of carrying out numerical computations in this work, the fermionic degrees of freedom in this Hamiltonian are mapped onto spin operators through the Jordan-Wigner transformation. Details of the above procedure can be found in Ref. [41]. Implementation of this strategy turns the Hamiltonian operator of the continuum Thirring model, $H_{{\mathrm{Th}}}$, into that of the XXZ spin chain coupled to both uniform and staggered magnetic fields, up to a scaling factor,
$H(\tilde{m}_{0}a,\Delta)=$ $-\frac{1}{2}\sum_{n}^{N-2}\left(S_{n}^{+}S_{n+1}^{-}+S_{n+1}^{+}S_{n}^{-}\right)$
$+$ $a\tilde{m}_{0}\sum_{n}^{N-1}\left(-1\right)^{n}\left(S_{n}^{z}+\frac{1}{2}\right)$
$+$ $\Delta(g)\sum_{n}^{N-1}\left(S_{n}^{z}+\frac{1}{2}\right)\ \left(S_{n+1}^{z}+\frac{1}{2}\right)\,,$ (2)
@@
where $a$ is the lattice spacing, $N$ is the total number of lattice sites, $S_{n}^{\pm}=S_{n}^{x}\pm iS_{n}^{y}$ and $S_{n}^{z}$ are the spin matrices ($S^{i}_{n}=\sigma^{i}/2$ with $\sigma^{i}$ being the Pauli matrices) at the $n$th site, and $[S^{i}_{n},S^{j}_{m}]_{n\not=m}=0$. Here $\tilde{m}_{0}=m_{0}/\nu(g)$ with $m_{0}$ being the bare counterpart of the mass parameter, $m$, in Eq. (1). The functions $\nu(g)$ and $\Delta(g)$ are the lattice version of wave-function renormalization and the four-fermion coupling [56],
$\nu(g)=$ $\left(\frac{\pi-g}{\pi}\right)/\sin\left(\frac{\pi-g}{2}\right)\,,$
$\Delta(g)=\cos\left(\frac{\pi-g}{2}\right)\,.$ (3)
@@
Notice that the physical Hamiltonian is obtained by rescaling the dimensionless one in Eq. (II.1) by a factor $\nu(g)/a$. Throughout the rest of the paper, we will work with the lattice Hamiltonian  (II.1) and correspondingly define dimensionless time and inverse-temperature parameters, which involve rescaling the physical (dimensionful) magnitudes by $\nu(g)/a$. Additionally, we restrict our consideration to the sector with total spin (magnetization) zero [41] and find the ground state by adding the penalty term (see Appendix  A).
@@ II.2 Matrix product state
which can then be approximated as a MPS with bond dimension $D$,
@@
We search for DQPTs in the real-time evolution of Eq. (II.1). For this purpose, it is desirable to perform calculations directly in the thermodynamic limit. This can be accomplished by employing the uniform MPS (uMPS) technique, which allows for the description of the state of an infinite spin chain with translation invariance using one bulk tensor and two appropriate boundary tensors [57, 58]. The uniform MPS can be represented as
@@
where $A$ is the unit cell of the uMPS. Due to the staggered term in Eq. (II.1), the system is translationally invariant with a period of two. Thus, we choose the tensor $A$ to represent two contiguous sites of the chain, and its physical dimension becomes $d=4$.
III Equilibrium phase structure of the 1+1-dimensional Thirring model
In this section, we review our study of the zero-temperature phase structure of the Thirring model using MPS [41]. An analysis of renormalization group (RG) equations of the dual sine-Gordon theory [59, 60, 61] points to the existence of two phases in the Thirring model. The massless theory is a conformal field theory at all couplings $\Delta(g)$. RG flows in the mass-coupling plane reveal that the line $m=0$ is a fixed line under RG transformations. However, it is stable/unstable for couplings $\Delta(g)$ below/above certain $\Delta(g_{\star})_{m=0}=-\pi/2$. At $m>0$, the nonzero fermion mass is an irrelevant operator for $\Delta(g)<\Delta(g_{\star})_{m}$ but becomes relevant when $\Delta(g)>\Delta(g_{\star})_{m}$, with the transition coupling $\Delta(g_{\star})_{m}$ being mass dependent. This gives rise to two equilibrium phases:
     (i)  $\Delta(g)<\Delta(g_{\star})_{m}$ : gapless (critical) phase
     (ii)  $\Delta(g)>\Delta(g_{\star})_{m}$ : gapped (massive) phase,
@@ III Equilibrium phase structure of the 1+1-dimensional Thirring model
$C_{\rm string}(x)=$ $\frac{1}{N_{x}}\sum_{n}\langle S^{z}(n)S^{z}(n+1)\cdots$
$\cdots S^{z}(n+x-1)S^{z}(n+x)\rangle,$ (8)
which is a string of $S^{z}$ operators. We note that Ref. [41] erroneously states this string correlator as the fermion-antifermion one, containing $\langle S^{+}(n)S^{z}(n+1)\ldots S^{z}(n+x-1)S^{-}(n+x)\rangle$ terms. At large-enough distances, the correlator $C_{\rm string}(x)$ exhibits a power-law decay in the gapless phase, given by $C_{\rm string}(x)\propto\beta x^{-\alpha}$, and a power-exponential decay in the gapped phase, $C_{\rm string}(x)\propto Bx^{-\eta}A^{x}$. The crucial distinction that makes the string correlator particularly useful for determining $\Delta(g_{\star})_{m}$ is that it decays to a constant, $C$, in the gapped phase. In practice, one can analyze both phases based on the power-exponential fitting ansatz:
@@
with $A=1$ found in fits in the gapless phase. Thus, the phase transition when going from the gapless into the gapped phase is manifested by $A$ dropping below 1 and, in the case of the string correlator, by $C>0$. The onset of the nonzero constant is numerically more robust than the smooth drop of $A$ below 1, accounting for the practical usefulness of the string correlator. As we will see below, this is true also at nonzero temperatures and it will allow us to identify relations between thermal phases of the Thirring model and DQPTs.
As a summary of our findings for the zero-temperature phase structure, we show Fig. 1. The color-coded regions correspond to the equilibrium phase on the coupling-mass plane through the correlator analyses from our previous work [41]. In particular, we fitted Eq. (9) and extracted values of the constant $C$. Since the results depend, in general, on the fitting interval, we adopted a systematic procedure to assess the systematic uncertainty of $C$ (dominating with respect to other uncertainties, coming from the finite bond dimension and the finite volume). For details of this procedure, we refer again to Ref. [41]. Depending on the value of $C$, we identified three situations. With $C<0.001$ or $C>0.01$, one unambiguously concludes either the critical phase (purple region) or the gapped phase (orange region), respectively. If $C\in[0.001,0.01]$ (gray region), then the systematic uncertainty does not allow us to draw conclusions about a (non)vanishing value of $C$. This region of ambiguous values of $C$ is indicated by gray shading and the BKT transition must occur there. We note the onset a nonzero value of $C$ is very smooth at small fermion masses, while it becomes rather sharp with an increasing mass. At very small masses, the implied transition point is consistent with the RG flow analysis [$\Delta(g)\approx-\pi/2$] and at large masses, it moves towards more negative values.